A homebuyer borrows 400,000 to be repaid over a 20 year period with level monthly payments beginning one month after the loan is made. The interest rate on the loan is a nominal annual rate of 12% convertible monthly. Find using mathematical formulas:

a. the total principal paid on the loan over the first 15 years

b. the total interest paid on the loan over the first 15 years
Answer #1

monthly payment or PMT formula

=PV/((1-(1+(r/m))^(-n*m))/(r/m))

=400000/((1-(1+(12%/12))^(-20*12))/(12%/12))

=4404.3445

a.

value of the loan at the end of 15 years=PMT*((1-(1+(r/m))^(-n*m))/(r/m))

=4404.3445*((1-(1+(12%/12))^(-5*12))/(12%/12))

=197997.48

the total principal paid on the loan over the first 15 years=loan amount-value of the loan at the end of 15 years

=400000-197997.48

=202002.52

b. the total interest paid on the loan over the first 15 years

=amount payment in 15 years-the total principal paid on the loan over the first 15 years

=(4404.3445*15*12)-202002.52

=590779.50

the above is answer..